Find the solution to each system by the substitution method. 2x - 5y = -1 and x - 3y = -3
Answer by drk(1908)

Here are the equations:
(i) 2x - 5y = -1
(ii) x - 3y = -3
Since we must use substitution, we must solve one equation for a variable. Let's solve (ii) for x, it is easiest.
(iii) x = 3y - 3.
Substitute (iii) into (i) to get
(iv) 2(3y-3) - 5y = -1.
Solve for y.
(v) 6y - 6 - 5y = -1
y = 5.
Solve for x.
x = 12.
(12, 5)